# Adaptive Noise Resilient Keyword Spotting Using One-Shot Learning

### v1

@@ Abstract
This study proposes a lightweight approach for continuous real-world noise adaptation in KWS using transfer learning on the last fully-connected layer of an NN model. The adaptation requires only one-shot learning and a single training epoch, significantly reducing data storage and computation requirements with respect to the state-of-the-art. This paper provides a proof-of-concept approach, which is designed to be deployable on resource-constrained devices in the future.
@@ III-A Datasets
The clean and noisy audio signals, sampled at 16 kHz, are converted to the log-Mel time-frequency domain using 25 ms windows, a 10 ms hop size, and 64 Mel-filters spanning 50–7500 Hz. This preprocessing step reduces the signal length by a factor of 2.48 and can be performed using an external low-power analog front-end, making it suitable for resource-constrained devices [4].
@@ III-B Method
The method described below first generates two pretrained models for KWS classification, which are then used to adapt to specific noise conditions.
The proposed model architecture, presented in detail in Fig. 2, takes the log-Mel spectrogram as input to classify one of the twelve keyword classes. The architecture comprises five convolutional layers interspersed with batch normalization layers, followed by an average pooling layer that computes the mean across the entire image channel. This single-channel value is then processed by a fully-connected layer.
@@
To enable adaptation to continuous noise conditions, this study explores the use of few-shot transfer learning techniques. Only the last fully-connected layer is fine-tuned, keeping all convolutional layers frozen to minimize memory and computational requirements. This technique allows for adjusting the pretrained models to changing noise scenarios over time, improving their robustness in shifting environments. In the future, the pretrained model and the continuous noise adaptation are intended to be deployed on resource-constrained devices for real-time operation. Therefore, a lightweight implementation is key.
@@ III-C Experimental Setup
The GCS dataset with clean utterances is divided into train, validation, and test sets in an 80:10:10 ratio. Each set contains audio clips balanced by class, as specified in the original dataset. Model performance is quantified using the accuracy metric. Additionally, the audio clips are mixed with individual noise sources to create noisy train, validation, and test sets with SNRs ranging from -3 to 24 dB in 3 dB increments.
@@
In this experiment, the Adam optimizer and cross-entropy loss function are used for training, with an initial learning rate of . The learning rate decays by a factor of if no improvement is observed for 10 consecutive epochs. Training is conducted with a batch size of 16 and concludes after 50 epochs. Default values are retained for all other parameters.
The second experiment evaluates the domain adaptation of the pretrained baseline-model and the best-performing noise-aware-model to specific on-site noise sources by applying few-shot transfer learning. This experiment investigates the number of training samples and epochs required for adaptation, as well as the models’ robustness to changing SNR conditions after adaptation. The number of samples per class (shots) is evaluated between 1 and 5, varying the number of training epochs from 1 to 5 for each case. During adaptation, backpropagation computes the stochastic gradient descent with a learning rate of and cross-entropy loss function.
@@ IV-A Comparison of pretrained models
The baseline-model, trained with noiseless utterance samples, achieves an accuracy of 96.34% for classifying clean (noise-free) keyword utterances. In comparison, the noise-aware-model instances, trained with clean signals plus 20%, 40%, or 60% additional noisy signals, achieve similar accuracies of 96.26%, 96.42%, and 96.11%, respectively. This represents a decrease of less than 0.22% compared to the baseline-model. However, the accuracy on a clean test set decreases by 1.06% and 0.49% for the noise-aware-model instances trained with an additional 80% and 100% of noisy signals, respectively.
Figure 3 presents the mean accuracy and standard deviation (SD) of the noise-aware-model when tested with different noise sources (colored and indoor noise) at SNRs ranging from -3 to 24 dB in 3 dB increments. The model instance trained with 100% additional noisy data achieves the best performance (higher mean accuracy and lower SD) for ”White”, ”Pink”, and ”Babble” noise sources. The model instance trained with 60% additional noisy data achieved the second-best performance. This instance also exhibits the highest mean accuracy for ”Kitchen”, ”Office”, and ”Living” noise sources. Given its overall superior performance across the six noise sources and only a minimal reduction in clean utterance accuracy, the noise-aware-model instance trained with an extra 60% of noisy data is considered in the remainder of this paper for continuous noise adaptation.
@@ IV-B On-site noise adaption
Both pretrained models are adapted to a specific on-site noise condition (source and SNR value) using 1-shot learning and 1 epoch to ensure minimal storage and processing requirements. Figure 5 shows the results of adapting the pretrained models at minimal and maximal SNRs (-3 dB and 24 dB, represented by orange and green lines, respectively). The adapted models are tested across the entire SNR range to account for the non-stationary nature of dynamic real-world noise conditions. Minimal accuracy variations, ranging from -0.22% to 2.54%, are observed for adaptations with other SNRs not shown in Fig. 5. The results indicate that the pretrained models effectively adapt to new conditions, achieving superior performance and robustness across different SNRs when compared to the original models (blue lines in Fig. 5). The best-performing adapted model depends on the specific noise scenario. Models adapted at -3 dB achieve higher accuracy at low and middle range SNRs (18 dB) than those adapted at 24 dB. In contrast, at higher SNRs (18 dB), the adapted noise-aware-model attains the highest accuracy, regardless of whether it was trained at -3 or 24 dB. However, at higher SNRs, the pretrained models before adaptation already exhibit accuracies above 90% for all noise sources, raising questions regarding the necessity of adaptation. Therefore, the conditions for adapting a continual learning model should be carefully selected.
@@
For large memory and computational devices, increasing the number of training samples and epochs may yield improvements in adaptation. Figure 6 shows the performance variations as a function of the increasing number of epochs used for adapting the pretrained baseline-model. The model is trained for 1 to 5 epochs with 1- or 5-shot learning utilizing utterances contaminated with specific noise sources and SNRs. Model evaluation is done at the same noise source and SNR value used during training adaptation. The results show that the model adapts well to noise, increasing its performance when trained for only 1 epoch. Adaptation capabilities can still be improved by increasing the number of training epochs, particularly for highly contaminated signals with noise (SNRs 3 dB), and in some cases for values between 3 and 9 dB. The performance remains relatively constant for SNRs higher than 9 dB. Note that for clarity reasons, we have chosen not to show all SNR conditions in Fig. 6.
The effect of increasing the number of training samples from 1-shot to 5-shot learning presents similar accuracy when adaptations are done for 1 epoch. Performance differences between 1-shot and 5-shot continual learning are observed as the number of epochs increases, especially for SNRs below 3 dB. For the ”Car horn” noise source, there is no clear improvement when adding more data during adaptation, while for ”Street music”, 1-shot learning achieves higher accuracy. For the ”Dog bark” noise source at SNR6 dB, 5-shot learning performs better than 1-shot, particularly as the number of training epochs increases. However, it is important to note that 5-shot learning requires storing and processing five times more data per epoch compared to 1-shot learning. Similar patterns are observed in the noise-aware-model when evaluating performance by increasing the number of epochs and data used during adaptation.
@@ V Conclusion
A key requirement of real-time KWS systems operation is their ability to adapt to changing noise conditions. To this end, this study proposed a few-shot transfer learning approach for continuous noise adaptation of pretrained NNs. We developed two pretrained models in a first experiment, a baseline-model trained on noiseless utterances and a noise-aware-model trained on a mix of noiseless and noisy utterances contaminated with colored and general indoor noises. The continuous noise adaptation capabilities of both pretrained models were rigorously evaluated under three specific on-site noise sources at different SNRs.

### v2

@@ Abstract
This study proposes a lightweight approach for continuous real-world noise adaptation in KWS using transfer learning (TL) on the last fully-connected layer of an NN model. The adaptation requires only one-shot learning and a single training epoch, significantly reducing data storage and computation requirements with respect to the state-of-the-art. This paper provides a proof-of-concept approach, which is designed to be deployable on resource-constrained devices in the future.
@@ III-B Method
The method described below first generates two pretrained models for KWS classification, which are then used to adapt to specific noise conditions. The end-to-end classification of keywords using acoustic signals begins with the computation of the log-Mel spectrogram. In this study, we follow the same procedure as [4] to convert the acoustic signal to the time-frequency domain. The clean and noisy audio signals, sampled at 16 kHz, are converted to the log-Mel time-frequency representation using 25 ms windows, a 10 ms hop size, and 64 Mel-filters spanning 50–7500 Hz. This preprocessing step reduces the signal length by a factor of 2.48. As demonstrated in [4], this log-Mel spectrogram configuration can be computed using an external low-power analog front-end, significantly reducing the energy consumption of an end-to-end keyword spotting system on resource-constrained devices.
The proposed model architecture, presented in detail in Fig. 2, takes the log-Mel spectrogram as input to classify one of the twelve keyword classes. Inspired by the CNN model used in [4], our architecture comprises five convolutional layers interspersed with batch normalization layers, followed by an average pooling layer that computes the mean across the entire image channel. This single-channel value is then processed by a fully-connected layer.
@@
To enable adaptation to continuous noise conditions, this study explores the use of few-shot TL techniques. Only the last fully-connected layer is fine-tuned, keeping all convolutional layers frozen to minimize memory and computational requirements. This technique allows for adjusting the pretrained models to changing noise scenarios over time, improving their robustness in shifting environments. In the future, the pretrained model and the continuous noise adaptation are intended to be deployed on resource-constrained devices for real-time operation. Therefore, a lightweight implementation is key.
@@ III-C Experimental Setup
The GCS dataset with clean utterances is divided into train, validation, and test sets in an 80:10:10 ratio. Each set contains audio clips balanced by class, as specified in the original dataset. Additionally, the audio clips are mixed with individual noise sources to create noisy train, validation, and test sets with SNRs ranging from -3 to 24 dB in 3 dB increments. Model performance is quantified using the accuracy metric.
@@
In this experiment, the Adam optimizer and cross-entropy loss function are used for training, with an initial learning rate of . The learning rate decays by a factor of if no improvement is observed for 10 consecutive epochs. Training is conducted with a batch size of 16 and concludes after 50 epochs. Default values are retained for all other parameters. The experiment is repeated 25 times using different model weight initializations to obtain reliable results.
The second experiment evaluates the domain adaptation of the pretrained baseline-model and the best-performing noise-aware-model to specific on-site noise sources by applying few-shot TL. This experiment investigates the number of training samples and epochs required for adaptation, as well as the models’ robustness to changing SNR conditions after adaptation. The number of samples per class (shots) is evaluated between 1 and 5, varying the number of training epochs from 1 to 5 for each case. During adaptation, backpropagation computes the stochastic gradient descent with a learning rate of and cross-entropy loss function.
@@ IV-A Comparison of pretrained models
In the initial experiment, both the baseline-model and each of the five instances of the noise-aware-model are trained 25 times with different weight initializations. The average accuracy over these 25 repetitions is used to evaluate the model’s performance. The baseline-model, trained with noiseless utterance samples, achieves a mean accuracy of 96.27% for classifying clean (noise-free) keyword utterances. In comparison, the noise-aware-model instances, trained with clean signals plus 20%, 40%, 60%, 80%, and 100% additional noisy signals, achieve a declining accuracies of 96.04%, 95.98%, 95.95%, 95.81%, and 95.73%, respectively. This represents a decrease of less than 0.54% compared to the baseline-model (p0.001 in all comparisons; Wilcoxon signed-rank test [28]).
Figure 3 presents the mean accuracy of the noise-aware-model instances when tested with different noise sources (colored and indoor noise) at SNRs ranging from -3 to 24 dB in 3 dB increments. While no single noise level added during training yields universally optimal performance across all noise sources, the 100% noise-trained noise-aware-model demonstrates superior accuracy for 3 of 6 noise sources and the highest overall mean accuracy. Statistical validation confirms this observation, with Friedman tests indicating significant performance variation between models (=1126.856, p0.001) and post-hoc Wilcoxon signed-rank tests (Holm-corrected) showing the 100% noise variant significantly outperforming all alternatives (all p0.001). Consequently, the noise-aware-model instance trained with 100% noise is selected as the optimal noise-aware-model for subsequent continuous noise adaptation experiments.
@@ IV-B On-site noise adaption
Both pretrained models are adapted to a specific on-site noise condition (source and SNR value) using 1-shot learning and 1 epoch to ensure minimal storage and processing requirements. Figure 5 shows the results of adapting the pretrained models at minimal and maximal SNRs (-3 and 24 dB, represented by orange and green lines, respectively). The adapted models are tested across the entire SNR range to account for the non-stationary nature of dynamic real-world noise conditions. For models adapted to intermediate SNR values (-3 dBSNR24 dB) not shown in Fig. 5, the accuracy curves consistently fall within the range delineated by the -3 and 24 dB curves. The results indicate that the pretrained models effectively adapt to new conditions, achieving superior performance and robustness across different SNRs when compared to the original models (blue lines in Fig. 5). The best-performing adapted model depends on the specific noise scenario. Models adapted at -3 dB achieve higher accuracy at low and middle range (SNR18 dB) than those adapted at 24 dB. In contrast, at SNR18 dB, the adapted noise-aware-model attains the highest accuracy, regardless of whether it was trained at -3 or 24 dB. However, at higher SNRs, the pretrained models before adaptation already exhibit accuracies above 90% for all noise sources, raising questions regarding the necessity of adaptation. Therefore, the conditions for adapting a continual learning model should be carefully selected.
@@
For large memory and computational devices, increasing the number of training samples and epochs may yield improvements in adaptation. Figure 6 shows the performance variations as a function of the increasing number of epochs used for adapting the pretrained baseline-model. The model is trained for 1 to 5 epochs with 1- or 5-shot learning utilizing utterances contaminated with specific noise sources and SNRs. Model evaluation is done at the same noise source and SNR value used during training adaptation. The results show that the model adapts well to noise, increasing its performance when trained for only 1 epoch. Adaptation capabilities can still be improved by increasing the number of training epochs, particularly for highly contaminated signals with noise (SNR3 dB), and in some cases for values between 3 and 9 dB. The performance remains relatively constant for SNRs higher than 9 dB. Note that for clarity reasons, we have chosen not to show all SNR conditions in Fig. 6.
The effect of increasing the number of training samples from 1-shot to 5-shot learning presents similar accuracy when adaptations are done for 1 epoch. Performance differences between 1-shot and 5-shot continual learning are observed as the number of epochs increases, especially for SNR3 dB. For the ”Car horn” noise source, there is no clear improvement when adding more data during adaptation, while for ”Street music”, 1-shot learning achieves higher accuracy. For the ”Dog bark” noise source at SNR6 dB, 5-shot learning performs better than 1-shot, particularly as the number of training epochs increases. However, it is important to note that 5-shot learning requires storing and processing five times more data per epoch compared to 1-shot learning. Similar patterns are observed in the noise-aware-model when evaluating performance by increasing the number of epochs and data used during adaptation.
@@ V Conclusion
A key requirement of real-time KWS systems operation is their ability to adapt to changing noise conditions. To this end, this study proposed a few-shot TL approach for continuous noise adaptation of pretrained NNs. We developed two pretrained models in a first experiment, a baseline-model trained on noiseless utterances and a noise-aware-model trained on a mix of noiseless and noisy utterances contaminated with colored and general indoor noises. The continuous noise adaptation capabilities of both pretrained models were rigorously evaluated under three specific on-site noise sources at different SNRs.